OpenStudy (anonymous):

Solve for n and select the correct answer. 18n + 29 = -9(4 – 2n)

OpenStudy (anonymous):

are u sure its a - inside the brackets? bcoz if it is then 18n will cancel 18n and we can't solve for n!

OpenStudy (anonymous):

18n+29=-36+18n 18n-18n=-36-29

OpenStudy (anonymous):

then theres no solution right?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

then?? what would it be

OpenStudy (anonymous):

is it all real solutions? or -36 or 29

OpenStudy (anonymous):

i was just trying to show u that 18n-18n=-36-29 0=-65 which means there's no solution for n